OpenStudy (rebecca1588):

Cheryl has 16 cupcakes made. She can make 48 cupcakes an hour. A customer asks for 160 cupcakes. How many hours will it take Cheryl to make enough cupcakes for the customer? I________I hours 3 Hours?

OpenStudy (anonymous):

is \(3\times 48+16=160\)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

You are right, it equals 3 hrs

OpenStudy (netlopes1):

48/h or 144/3h plus 16 = 160.... ok 3h

OpenStudy (anonymous):

yes it is so you win

OpenStudy (netlopes1):

Really, you are right.

OpenStudy (rebecca1588):

Thanks people!!! :D